OpenStudy (anonymous):

Where does the line that passes through the points P=(3,-1,2) and Q=(5,8,4) pierce the X-Y plane?

OpenStudy (amistre64):

define the vector between them first

OpenStudy (amistre64):

once you have the vector, you can create the equation to define every point on the line x = xo + at y = yo + bt z = zo + ct and since the xy plane is the z=0 plane; use the z equation to define t for z=0; then use that to determine x and y

OpenStudy (amistre64):

or, we can continue the shuffiling generally 0 = zo + ct ; when t=-zo/c ; therefore the point will be (xo + a(-zo/c), yo + b(-zo/c), zo + c(-zo/c))

OpenStudy (anonymous):

Ok I think am getting some idea , but I actually jxt saw the question and cudnt ans it!

OpenStudy (amistre64):

aslong as we know how to define a vector between 2 points, the general setup of the point just needs filled in by the <a,b,c> vector parts:)

OpenStudy (amistre64):

and of course the xo,yo,zo are one of the points that had been given

OpenStudy (anonymous):

Oh Ok! I will try to solve it then! Anyway thanks for your help!!!!!!!!!!!